OpenStudy (anonymous):

what is the answer too 3x(x^2-7x+3)-4x(2+19x-2x^2)

OpenStudy (anonymous):

Expand and collect like-terms: 3x^3-21x^2+9x-8x-76x^2+8x^3 =11x^3-97x^2+x = x(11x^2-97x+1)

OpenStudy (anonymous):

=3x^3-21x^2+9x-8x-76x^2+8x^3 =11x^3-97x^2+x